OpenStudy (anonymous):

the sum of two numbers is 1660 and HCF is 20 find the numbers.

OpenStudy (anonymous):

Let the numbers be 20a and 20b because 20 is the HCF, so HCF is the highest common factor by which both the numbers can be divided.. Now, 20a + 20b = 1660 (Given in question) a + b = 83 Now, look for co-primes numbers that make the sum 83..(Co-primes are the pair of prime numbers whose HCF is 1 like (3,5), (5,8), means you can not divide two numbers by One single number or having no common factor between them)... You will have your answer...

OpenStudy (anonymous):

lol which is???

OpenStudy (anonymous):

Sorry, it is not neccessary that numbers will be prime.. Co-prime are the pair of numbers (not neccessarily prime) whose HCF is 1...

OpenStudy (anonymous):

There are a couple coprime pairs. I was wondering if this question was supposed to have only one answer or not.

OpenStudy (anonymous):

r u sure the question is right? because 20*20+20*20=800 which is the highest possible sum you cud get with a HCF of 20

OpenStudy (anonymous):

jeomath what are they???

OpenStudy (anonymous):

I also, think there is something went wrong in question given...

OpenStudy (anonymous):

(1, 82) (2,81) (3,80), those are some coprime pairs.

OpenStudy (anonymous):

the question is wrong

OpenStudy (anonymous):

just to confirm, HCF is the same as GCD right?

OpenStudy (anonymous):

But, HCF of every one is not 20... So, I think question is wrong...

OpenStudy (anonymous):

dere are no two numbers with an HCF of 20 dat can add up to 1660, because 20*20+20*20= 800 cant go over 800 without a higher HCF

OpenStudy (anonymous):

1600 + 60 = 1660, and the HCF of 1600 and 60 is 20.

OpenStudy (anonymous):

Ya Highest Common Factor And Greatest Common Divisor are the one and the same thing...

OpenStudy (anonymous):

40 is a factor of 1600

OpenStudy (anonymous):

ohh true sorry

OpenStudy (anonymous):

i was wrong

OpenStudy (anonymous):

20+ 1640 = 1660, HCF is again 20.

OpenStudy (anonymous):

40+ 1620 = 1660, again HCF is 20, theres not a unique answer.

OpenStudy (anonymous):

Oh sorry. (80, 3) is the perfect pair which makes the sum = 83... Now the numbers were 20a = 20*8 = 1600 and second number is 20b = 20*3 = 60 is the answer...

OpenStudy (anonymous):

But as (81,2) is also a pair which satisfies the same condition, so this question is not having a unique or particular answer...

OpenStudy (saranya):

the nos. can be written as 20a and 20b so here is it , 20a + 20b = 1660 => a + b = 83 a and b can take any value like if a=1,b=82 ie nos are 20 and 1640 but a and b should be co-prime